# Hyperparameter Optimization in Machine Learning

### v2

@@ 4.2.2 Gaussian processes
Instead of worrying about feature maps and weights, we can directly work with a kernel function, a notion which gives rise to Gaussian processes (GP) [williams2006gaussian, murphy2012machine]. In fact, the Bayesian linear model converges to a GP in the limit of infinitely many features, if the prior distribution on the weights is scaled appropriately [neal1996bayesian].
A Gaussian process is a Gaussian distribution over functions $g(\lambda)$, fully specified by a mean function $\mu(\lambda)$ and a covariance function $\kappa(\lambda,\lambda^{\prime})$, or kernel function. It has the property that for any finite set $\lambda_{1},\dots,\lambda_{n}\in\Lambda$, the joint distribution of the vector $[g(\lambda_{1}),\dots,g(\lambda_{n})]$ is multivariate Gaussian with mean $[\mu(\lambda_{i})]_{i=1,\dots,n}$ and covariance matrix $[\kappa(\lambda_{i},\lambda_{j})]_{i,j=1,\dots,n}$. Our linear model derivation above gives rise to a zero-mean Gaussian process (since the prior over weights (4.8) has mean zero) with covariance function $\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta$, where $\delta(\lambda,\lambda^{\prime})=1$ if $\lambda=\lambda^{\prime}$ and 00 otherwise. In short, we can write
$\hat{f}\sim\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta).$
The surrogate model update equation $\pi_{2}$ (see (4.3)) may then be written as
$\hat{f}_{k+1}=\pi_{2}\left(\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k},\xi_{2}\right)\sim\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta\,|\,X_{k},Y_{k}),$ (4.12)
which expresses the update as the conditioning of the Gaussian process on the observations $X_{k}$, $Y_{k}$. While this posterior is in principle an infinite-dimensional object, we only ever need its marginal distributions over single points $\lambda_{*}$ (or over small sets, such as pairs) in order to compute AFs and make sampling decisions. These marginals are Gaussian with mean and (co)variance given by simple linear algebra expressions (4.11).
@@
Gaussian processes also come with a number of drawbacks that may limit their applicability to real-world scenarios. Most importantly, computing the posterior for $k$ datapoints costs $\mathcal{O}(k^{3})$, as can be seen in (4.11). If we are able to sample $f$ a few thousand times or more, GP-based BO starts to become intractable. It is also difficult to design kernel functions for high-dimensional $\Lambda$, or if the search space contains many integer, categorical, or conditional hyperparameters. In such situations, BO with other surrogate models can perform better, as shown in a comparative study by eggensperger_towards_2013.
snoek_practical_2012 tackle the first problem, proposing an extension of the expected improvement that accounts for pending executions, and take on the scalability issues in snoek_scalable_2015 by replacing the GP with a Bayesian neural network. hutter_sequential_2011, bergstra2011algorithms and bergstra_making_2013 focus both on scalability and applicability to discrete and conditional search spaces by using random forests and tree Parzen estimators (see also Section 4.2.5). More recent advances in the field [springenberg2016bayesian, wu2017bayesian, klein2017fast] focus on replacing GPs with more practical and scalable approaches, exploiting additional information besides function evaluations (thus “dropping” the black-box assumption for $f$). The additional information could be, for example, the performance of $A$ on a small subsets of $D$ [klein2017fast] or (possibly noisy) gradients [wu2017bayesian]. Finally, falkner2018bohb propose hybridizing BO with bandit-based methods (discussed in the next section) to improve performance in the first stage of HPO, when the surrogate model may still be rather unreliable.
@@ 4.2.3 Acquisition functions
Suppose we fitted our surrogate model $\hat{f}_{k}$ to the current set of evaluations $\mathcal{D}=\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k}$. To decide which point to evaluate next, BO optimizes an acquisition function $\alpha(\lambda|\hat{f}_{k})$. In other words, our suggestion mapping $\pi_{1}$ becomes
$\lambda_{k+1}=\pi_{1}(\hat{f}_{k},\nu_{2})=\arg\max_{\lambda\in\Lambda}\,\alpha(\lambda|{\hat{f}_{k}}).$ (4.13)
In this section, we restrict our attention to Gaussian process surrogate models, even though most AFs discussed here can be applied to other models just as well. As we will see, most common AFs depend on the posterior distribution of $\hat{f}_{k}$ only through its marginal mean and variance
$\mu_{k}(\sigma)=\mathbb{E}\left[\hat{f}_{k}(\lambda)\right],\quad\sigma_{k}(\lambda)^{2}=\mathbb{V}\left[\hat{f}_{k}(\lambda)\right],$
given by (4.11). Here, $\mathbb{E}$ and $\mathbb{V}$ denotes expectation and variance over the posterior distribution of $\hat{f}_{k}(\lambda)$ respectively.
A useful AF has to guide the choice of the next candidate so as to gain as much new information about high-performing hyperparameters as possible, given a limited sample $\mathcal{D}$ obtained so far. In particular, it needs to balance the trade-off between exploration (i.e., looking away from points in $\mathcal{D}$ in order to increase coverage) and exploitation (i.e., refining the search by probing near current optima in $\mathcal{D}$). We can gain intuition by considering two extremes. First, we could sample at our current best guess of where the minimum should be:
$\alpha_{\text{exploit}}(\lambda|\hat{f}_{k})=-\mu_{k}(\lambda).$
@@
$\alpha_{\text{explore}}(\lambda|\hat{f}_{k})=\sigma_{k}(\lambda).$
@@
$\alpha_{\text{LCB}}(\lambda|\hat{f}_{k})=-\mu_{k}(\lambda)+\eta\sigma_{k}(\lambda),\quad\eta>0,$
is a competitive and frequently used AF, called lower-confidence bound (LCB) [Srinivas2009].
@@
$\alpha_{\text{EI}}(\lambda|\beta,\hat{f}_{k})\doteq\operatorname{\mathds{E}}\left[\max\left\{\beta-\hat{f}_{k}(\lambda),0\right\}\right].$
Here, $\beta\in\mathbb{R}$ is a baseline, which is typically chosen as the minimum value of $f$ observed so far, $\beta=f(\hat{\lambda}_{k})$. In contrast to LCB, EI does not come with a free parameter to be set. In addition, it can be shown to have a simple closed-form expression in terms of $\mu_{k}(\lambda)$ and $\sigma_{k}(\lambda)$:
$\alpha_{\text{EI}}(\lambda|\beta,\hat{f}_{k})=\sigma_{k}(\lambda)\nu(\lambda)\Phi\left(\nu(\lambda)\right)+\sigma_{k}(\lambda)\phi\left(\nu(\lambda)\right)$ (4.14)
@@
$\nu(\lambda)\doteq\frac{\beta-\mu_{k}(\lambda)}{\sigma_{k}(\lambda)}$
@@
Equations (4.13) where $\alpha$ is the expected improvement and (4.12) specify a particular simple instantiation of a GP-based BO algorithm that may be used for HPO. As an illustrative example, two steps of the Bayesian optimization algorithm using EI as an AF are shown in Figure 4.1.
Figure 4.1: Left: Suppose we have measured the response function (plotted in red) on a small set of points. Fitting a Gaussian process on these points yields a mean function (plotted in black) with and a variance function ($\pm$ one standard deviation is filled in gray). If $\lambda$ falls between 7 and 11, the uncertainty is very high, which yields the right bump in the expected improvement (plotted in green), allowing us to explore a region where potentially a smallest response might be found. The left bump of the expected improvement falls in an exploitation region, which in this example looses against exploration. Hence, the next candidate is $\lambda=8.5$. Right: once the response is measured at the new candidate, the Gaussian process is trained again. Now uncertainty shrinks significantly and the new expected improvement has just one bump around 2.7, which is already very close to the minimum of the true response. This example takes inspiration from Figure 11 in [Jones1998:Efficient].
@@
An earlier and simpler variant of EI is the probability of improvement [Kushner1964], given by
$\alpha_{\text{PI}}(\lambda|\beta,\hat{f}_{k})\doteq P\left[\hat{f}_{k}(\lambda)<\beta\right]=\Phi(\nu(\lambda)).$ (4.15)
Another popular AF is known as Thompson sampling (TS) [Thompson1933, Hoffman2014]. The idea is to draw a sample path from our surrogate model and to search for its minimum point. Exact TS is intractable for a GP surrogate model, sample paths of which do not have a finite representation. In practice, we can draw a joint sample over a finite set of points333 Which costs $\mathcal{O}(b^{3})$ time and $\mathcal{O}(b^{2})$ memory, if $b$ is the size of the set. [unify notation please] or approximate the GP posterior by a Bayesian linear model [Hernandez-Lobato2014].
Other relevant families of AFs rely on information-theoretic criteria [Villemonteix2009, Hennig2012, Hernandez-Lobato2014, wang2017max]. While these have a sound theoretical foundation and have been shown to be more sample efficient, they typically come with higher computational cost and are more complex to implement than standard options such as EI or LCB. Given that the choice of the best AF can depend on the target function to optimize, one can also resort to a multi-armed bandit strategies to pick the most suitable AF from a portfolio of choices over the course of the optimization [hoffman2011portfolio, shahriari:2016:taking_human_out].
Optimizing acquisition function in practice
Techniques for maximizing AFs include grid search, where the AF is evaluated at a predefined set of points, and random search, which evaluates the AF at randomly selected points. More sophisticated methods involve the use of gradient-based optimization, which leverages the gradient of the AF to find local maxima more efficiently.
@@ 4.2.4 Aleatoric and epistemic uncertainty
In (4.4) we have introduced a noise term $\epsilon$ in order to account for the stochastic nature of the learning algorithm itself. This derives from the internal randomness of the training algorithm, (for example due to initial weights when training a deep network). In general, the amount of this aleatoric uncertainty may depend on the current hyperparameters and should not be confused with the epistemic uncertainty which arises because we may have not (yet) evaluated a large enough set of hyperparameter vectors. The latter uncertainty is crucial to allow explorations in regions not yet assessed, while the former should not mislead us to acquire points where uncertainty cannot be reduced.
A heteroscedastic approach can help disentangle these types of uncertainties, as proposed by [griffiths:2021]. In their approach, distinct Gaussian processes are trained aiming to predict the unknown response and the noise separately.
@@ 4.2.5 Tree Parzen estimator
The Tree Parzen Estimator (TPE) is a probabilistic model to efficiently explore complex search spaces. It models the objective function as a pair of conditional probability distributions: one for configurations that yield better results than a threshold and another for those that perform worse. By iteratively selecting hyperparameter values that maximize the expected improvement, TPE focuses the search on promising regions. This approach is particularly well-suited for high-dimensional, non-convex optimization problems and balances exploration and exploitation to accelerate convergence toward optimal configurations. While Gaussian process based BO methods model the distribution $p(y|\lambda)$ of observations $y$ given configurations $\lambda$, the Tree Parzen Estimator (TPE) [bergstra2011algorithms, bergstra_making_2013] fits two conditional densities over $\Lambda$:
$l_{\beta}(\lambda)=p(\lambda|y<\beta),\quad g_{\beta}(\lambda)=p(\lambda|y\geq\beta).$ (4.16)
Here, the baseline $\beta$ is set to a lower $\gamma$-quantile of the observed targets $\{y_{i}\}$ (e.g., $\gamma=0.15$), giving rise to a $(\gamma,1-\gamma)$ split of the data, on which $l_{\beta}(\lambda)$, $g_{\beta}(\lambda)$ are fit respectively. While TPE uses a tree of Parzen estimators for conditional search spaces, later work made use of kernel density estimators [falkner2018bohb]. The density ratio $l_{\beta}(\lambda)/g_{\beta}(\lambda)$ is a strictly increasing function of the expected improvement444 However, note that we cannot set $\beta$ to the minimum value observed so far, as in the GP-BO case, since this implies $\gamma=0$ and no data to fit $l_{\beta}(\lambda)$ with. AF with baseline $\beta$, and TPE maximizes this AF approximately by sampling a set of configurations from $l_{\beta}(\lambda)$ and suggesting the one with the largest ratio $l_{\beta}(\lambda)/g_{\beta}(\lambda)$. TPE is conceptually simple, easy to run with parallel evaluations, and can naturally deal with conditional search spaces. Subsequent work has shown that density ratios can be estimated by way of probabilistic binary classification, which is simpler and better supported by ML methodology than density estimation [Tiao:21].
4.3 Other approaches to model-based HPO
Not all model-based methods follow a Bayesian optimization framework. ilievski2017efficient use a deterministic radial basis function model for $\hat{f}$ which they interpolate on observations of $f$, while hazan2017hyperparameter focus on Boolean hyperparameters and use a sparse Fourier basis model for $\hat{f}$, in conjunction with heuristics to estimate the most sensitive hyperparameter entries. They then use a base global optimizer to perform HPO on a restriction of the original hyperparameter space. The works by lorraine2018stochastic and mackay2019self, linked to gradient-based HPO may also be interpreted as an attempt to learn a surrogate model of the entire learning algorithm (applied to a specific dataset), whereby the output space of $\hat{f}$ coincides with the hypothesis space of $A$ itself, that is $\hat{\mathcal{Y}}=\mathcal{H}$ in (4.1). We cover gradient-based HPO in Chapter 7.
Furthermore, not all methods follow the iterative schema of Eq. (4.3). Another possibility is to first learn $\hat{f}$ and then use it to suggest one or several hyperparameter configurations in a second stage, following a parallel scheme. Examples of this latter approach include [feurer2015initializing, fusi2018probabilistic]. These methods, however, assume the presence of multiple tasks and are closer in nature to meta or transfer learning settings.
4.4 Example: model-based vs random search
@@
We illustrate the difference between model-based and random search in a concrete example of tuning regularization parameters of an ML model to prevent overfitting. We tuned the L1 and L2 regularization terms, respectively alpha and lambda, of the XGBoost algorithm on the UCI direct marketing binary classification dataset from Kaggle. The goal was to minimize the AUC. Figure 4.2 presents a comparison between random search and BO (model-based), which models the objective function through a GP and uses the expected improvement as the acquisition function.

### v3

@@ 4.2.2 Gaussian Processes
Instead of working with explicit feature maps and weights, we can directly define a model through a kernel function, which leads to the framework of Gaussian processes (GPs) [williams2006gaussian, murphy2012machine]. In fact, the Bayesian linear model described above converges to a GP in the limit of infinitely many features, provided the prior over the weights is scaled appropriately [neal1996bayesian].
A Gaussian process (GP) is a distribution over functions whose finite-dimensional marginals are multivariate Gaussian distributions. A GP is fully specified by a mean function $\mu\mathrel{\mathop{\ordinarycolon}}\Lambda\to\mathds{R}$ and a covariance function $\kappa\mathrel{\mathop{\ordinarycolon}}\Lambda\times\Lambda\to\mathds{R}$, also called the kernel. That is,
$g\sim\mathcal{GP}(\mu,\kappa)$
means $g$ is a random function such that for any finite set of points $\lambda_{1},\dots,\lambda_{k}\in\Lambda$, the random vector $(g(\lambda_{1}),\dots,g(\lambda_{k}))$ follows the multivariate Gaussian distribution:
$(g(\lambda_{1}),\dots,g(\lambda_{k}))\sim\mathcal{N}\left(M,\Upsilon\right)$
where the mean vector is given by $M_{i}=\mu(\lambda_{i})$ and the covariance matrix is given by $\Upsilon_{ij}=\kappa(\lambda_{i},\lambda_{j})$.
In Gaussian process-based Bayesian optimization, the surrogate model is a random function $\hat{f}\sim\mathcal{GP}(\mu,\kappa)$ taking values in the space of functions $\Lambda\to\mathds{R}$. The update of the Gaussian process after observing $\mathcal{C}_{k}$ (the fitting step in Algorithm 1) corresponds to conditioning the prior on the dataset $(X_{k},Y_{k})$, and can be written as:
$\hat{f}\sim\mathcal{GP}(\mu,\kappa\mid X_{k},Y_{k})=\mathcal{GP}(\mu_{k},\kappa_{k}),$ (4.10)
where typically $\mu$ is taken to be 0. While the posterior GP is technically an infinite-dimensional object, in practice we only require its marginal distributions at single query points $\lambda_{*}$ (or small batches of points) — which are (multivariate) Gaussians — to compute acquisition functions and guide the search.
Our derivation of the Bayesian linear model in the previous section yields a zero-mean Gaussian process (since the prior over weights, Eq. (4.6), has mean zero) with covariance function
$\kappa(\lambda,\lambda^{\prime})=\kappa_{\Sigma}(\lambda,\lambda^{\prime})+\sigma_{\varepsilon}^{2}\delta(\lambda,\lambda^{\prime})=\varphi(\lambda)^{\intercal}\Sigma\varphi(\lambda^{\prime})+\sigma_{\varepsilon}^{2}\delta(\lambda,\lambda^{\prime}),$
where $\delta(\lambda,\lambda^{\prime})$ is the Dirac delta function, and the term $\sigma_{\varepsilon}^{2}\delta$ accounts for independent Gaussian observation noise. More specifically, the model derived in Eq. (4.9) corresponds to the posterior predictive distribution of the GP at a single query point $\lambda_{*}$, conditioned on the observed dataset $(X_{k},Y_{k})$. This marginal distribution over $\mathds{R}$ is a univariate Gaussian and provides the mean
$\mu_{k}(\lambda_{*})=k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}Y_{k}$
and variance
$\kappa_{k}(\lambda_{*},\lambda_{*})=\kappa_{\Sigma}(\lambda_{*},\lambda_{*})-k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}k_{*}$
required for the evaluation of acquisition functions.
@@
Gaussian processes also come with a number of drawbacks that may limit their applicability to real-world scenarios. Most importantly, computing the posterior for $k$ data points costs $\mathcal{O}(k^{3})$, due to the inversion of Eq. (4.9). If we are able to sample $f$ a few thousand times or more, GP-based BO starts to become intractable. It is also difficult to design kernel functions for high-dimensional $\Lambda$, or if the search space contains many integer, categorical, or conditional hyperparameters. In such situations, BO with other surrogate models can perform better, as shown in a comparative study by eggensperger_towards_2013.
snoek_practical_2012 tackle the first problem, proposing an extension of the expected improvement that accounts for pending executions. snoek_scalable_2015 tackles scalability issues by replacing the GP with a Bayesian neural network. hutter_sequential_2011, bergstra2011algorithms, bergstra_making_2013, perrone2018scalable, jenatton2017bayesian focus on scalability and applicability to discrete and conditional search spaces by using random forests and tree Parzen estimators, which we will discuss in Section 4.2.5). More recent advances in the field [springenberg2016bayesian, wu2017bayesian, klein2017fast] focus on replacing GPs with more practical and scalable approaches, exploiting additional information besides function evaluations (thus “dropping” the black-box assumption for $f$). The additional information could be, for example, the performance of $A$ on a small subsets of $D$ [klein2017fast] or (possibly noisy) gradients [wu2017bayesian]. Finally, falkner2018bohb propose hybridizing BO with bandit-based methods (discussed in the next chapter) to improve performance in the first stage of HPO, when the surrogate model may still be unreliable.
@@ 4.2.3 Acquisition Functions
We now come to the second central component of a Bayesian optimization method: the acquisition function (AF). Suppose we have fitted our surrogate model $\hat{f}_{k}$ to the current set of evaluations $\mathcal{C}_{k}=\{(\lambda_{j},f(\lambda_{j}))\}_{j=1}^{k}$. To decide which point to evaluate next, BO optimizes an acquisition function $\alpha(\lambda,\hat{f}_{k})$:
$\lambda_{k+1}=\arg\max_{\lambda\in\Lambda}\,\alpha(\lambda,\hat{f}_{k}).$ (4.11)
In this section, we primarily think of the surrogate model as a $\Lambda\to\mathds{R}$-valued random function modeled by a Gaussian process, as described above. However, most AFs discussed here can be applied to other models as well. As we will see, most common AFs depend on the posterior distribution of $\hat{f}_{k}$ only through its marginal mean and variance
$\mu_{k}(\lambda)=\mathbb{E}\left[\hat{f}_{k}(\lambda)\right],\quad\sigma_{k}(\lambda)^{2}=\mathbb{V}\left[\hat{f}_{k}(\lambda)\right],$
given by Eq. (4.9), where $\mathbb{E}$ and $\mathbb{V}$ denote expectation and variance, respectively.
A useful AF should guide the choice of the next candidate to gain as much new information about high-performing hyperparameters as possible, given the data $\mathcal{C}_{k}$ collected so far. In particular, it needs to balance the trade-off between exploration (i.e., looking away from points in $\mathcal{C}_{k}$ in order to increase coverage) and exploitation (i.e., refining the search by probing near current optima in $\mathcal{C}_{k}$). We can gain intuition by considering two extremes. First, we could sample at our current best guess of where the minimum should be:
$\alpha_{\text{exploit}}(\lambda\mid\hat{f}_{k})=-\mu_{k}(\lambda).$
@@
$\alpha_{\text{explore}}(\lambda,\hat{f}_{k})=\sigma_{k}(\lambda).$
@@
$\alpha_{\text{LCB}}(\lambda,\hat{f}_{k})=-\mu_{k}(\lambda)+\eta\sigma_{k}(\lambda),\quad\eta>0,$
is a competitive and frequently used AF, called lower-confidence bound (LCB) [Srinivas2009]. The choice of $\eta$ explicitly determines the trade-off between exploration and exploitation. Choosing $\eta$ depends on the application and, unfortunately, there is no good default rule of thumb to set the value.
@@
$\alpha_{\text{EI}}(\lambda,\hat{f}_{k},\beta)=\operatorname{\mathds{E}}\left[\max\left\{\beta-\hat{f}_{k}(\lambda),0\right\}\right].$
Here, $\beta\in\mathds{R}$ is a baseline, which is typically chosen as the minimum value of $f$ observed so far: $\beta=\min_{k\in[K]}f(\lambda_{k})$. In contrast to LCB, EI does not come with a free parameter to be set. EI makes an implicitly explore-exploit trade-off that has been reported to result in good performances in practice. A generalized version of EI replaces 0 by $\eta>0$, which adjusts this trade-off with a cut-off hyperparameter. In addition, it can be shown to have a simple closed-form expression in terms of $\mu_{k}(\lambda)$ and $\sigma_{k}(\lambda)$:
$\alpha_{\text{EI}}(\lambda,\hat{f}_{k},\beta)=\sigma_{k}(\lambda)\nu(\lambda)\Phi\left(\nu(\lambda)\right)+\sigma_{k}(\lambda)\phi\left(\nu(\lambda)\right)$ (4.12)
@@
$\nu(\lambda)=\frac{\beta-\mu_{k}(\lambda)}{\sigma_{k}(\lambda)}$
@@
Eq. (4.11), where $\alpha$ is the expected improvement, and Eq. (4.10) specify a particular simple instantiation of a GP-based BO algorithm that may be used for HPO. As an illustrative example, two steps of the Bayesian optimization algorithm using EI as the acquisition function are shown in Figure 4.1.
Figure 4.1: Left: Suppose we have measured the response function (plotted in red) on a small set of points. Fitting a Gaussian process on these points yields a mean function (plotted in black) with and a variance function ($\pm$ one standard deviation is filled in gray). If $\lambda$ falls between 7 and $1$, the uncertainty is very high, which yields the right bump in the expected improvement (plotted in green), allowing us to explore a region where potentially a smallest response might be found. The left bump of the expected improvement falls in an exploitation region, which in this example looses against exploration. Hence, the next candidate is $\lambda=8.5$. Right: once the response is measured at the new candidate, the Gaussian process is trained again. Now uncertainty shrinks significantly and the new expected improvement has just one bump around 2.7, which is already very close to the minimum of the true response. This example takes inspiration from Figure 11 in [Jones1998:Efficient].
@@
An earlier and simpler variant of EI is the probability of improvement [Kushner1964:New], given by
$\alpha_{\text{PI}}(\lambda,\hat{f}_{k},\beta)=\mathbb{P}\left[\hat{f}_{k}(\lambda)<\beta\right]=\Phi(\nu(\lambda)).$ (4.13)
Another popular AF is known as Thompson sampling (TS) [Thompson1933, Hoffman2014]. The idea is to draw a sample path from our surrogate model and to search for its minimum point. Exact TS is intractable for a GP surrogate model, sample paths of which do not have a finite representation. In practice, we can draw a joint sample over a finite set of points or approximate the GP posterior by a Bayesian linear model [Hernandez-Lobato2014].
Other relevant families of AFs rely on information-theoretic criteria [Villemonteix2009, Hennig2012, Hernandez-Lobato2014, wang2017max]. These typically are more complex to implement, come with higher computational cost, and require biased approximation, but have compelling theoretical foundations and often perform better than standard choices such as EI or LCB, achieving lower sample complexity. Given that the choice of the best AF can depend on the target function to optimize, one can also resort to a multi-armed bandit strategies to pick the most suitable AF from a portfolio of choices over the course of the optimization [hoffman2011portfolio, shahriari:2016:taking_human_out].
Optimizing acquisition functions in practice.
The optimization of acquisition functions can itself pose a challenge. As a result, Eq. (4.11) is rarely solved exactly. Techniques for maximizing AFs include grid, random, and quasi-random search, discussed in the previous chapter. Because surrogate models are cheap to evaluate, it is typically feasible to sample the acquisition function at many candidate points. This allows even naive search strategies to perform reasonably well in practice. More sophisticated approaches rely on gradient-based optimization, which uses the derivatives of the acquisition function to locate local maxima more efficiently. This is especially effective when the acquisition function is smooth and differentiable — such as in the case of GP-based posteriors where gradients can be computed analytically or via automatic differentiation. However, it is important to note that thorough optimization of the acquisition function is rarely essential. This is because the surrogate model is only an approximation of the true objective, and the acquisition function itself is a heuristic derived from it. Approximate maximization is typically sufficient to drive effective search progress for the overall HPO algorithm [Wilson2018].
@@ 4.2.4 Aleatoric and Epistemic Uncertainty
The noise term $\varepsilon$ introduced in Eq. (4.2) accounts for the inherent stochasticity of the learning algorithm—for example, due to random initialization when training a neural network. This kind of randomness gives rise to aleatoric uncertainty, which reflects intrinsic noise in the observations and may depend on the specific hyperparameter configuration.
However, as we saw in the previous sections, Bayesian optimization fundamentally relies on a surrogate model that not only predicts performance but also quantifies uncertainty. In particular, acquisition functions such as expected improvement and lower-confidence bound use the predictive variance of the surrogate model to guide exploration. This variance primarily captures epistemic uncertainty — uncertainty due to limited data. Crucially, this is the kind of uncertainty that can be reduced by collecting more observations, and that the acquisition function should leverage when choosing new points to evaluate.
Failing to distinguish between epistemic and aleatoric uncertainty can be problematic. If the surrogate model attributes too much of its predictive variance to aleatoric noise (which cannot be reduced), the acquisition function may over-prioritize regions where further evaluation is uninformative. Conversely, underestimating aleatoric noise can lead to overconfident predictions that neglect areas of genuine variability. To address this, heteroscedastic models—which explicitly account for input-dependent noise—can be employed. For example, griffiths:2021 train two distinct Gaussian processes aiming to predict the unknown response and the noise separately.
@@ 4.2.5 Tree Parzen Estimator
We now return to surrogate models and describe an implementation particularly well-suited for structured search spaces: the tree Parzen estimator (TPE). Unlike the Gaussian process framework, which models the conditional distribution of the response function given a configuration, TPE [bergstra2011algorithms, bergstra_making_2013] adopts a different approach. It models two conditional densities over configurations, given a performance threshold: one for configurations with good performance, and the other for those with poor performance. Formally, in the TPE framework the surrogate model $\hat{f}\mathrel{\mathop{\ordinarycolon}}\Lambda\times\mathds{R}\to\mathds{R}^{2}$ is defined as
$\hat{f}(\lambda,\beta)=\left(l_{\beta}(\lambda),\,g_{\beta}(\lambda)\right),$ (4.14)
where $l_{\beta}$ and $g_{\beta}$ are conditional density estimates over $\Lambda$. The threshold $\beta$ is typically set to the $\gamma$-quantile of the observed responses at iteration $k$, for some $\gamma\in(0,1)$, e.g., $\gamma=0.15$. This induces a $(\gamma,1-\gamma)$ partition of the data $\mathcal{C}_{k}$, where the set of good-performing configurations is defined as
$\{(\lambda,f(\lambda))\in\mathcal{C}_{k}\mid f(\lambda)<\beta\}\subset\mathcal{C}_{k},$
based on the available observations and the corresponding partition induced by $\gamma$. At each iteration $k$, the densities $l_{\beta}$ and $g_{\beta}$ are fitted separately to the two subsets of $\mathcal{C}_{k}$, so that
$l_{\beta}(\lambda)\approx p(\lambda\mid f(\lambda)<\beta),\quad g_{\beta}(\lambda)\approx p(\lambda\mid f(\lambda)\geq\beta).$
TPE uses as acquisition function the ratio of the two densities:
$\alpha_{\mathrm{TPE}}(\lambda,\hat{f}_{k},\beta)=\frac{l_{\beta}(\lambda)}{g_{\beta}(\lambda)},$
which assigns higher values to configurations that the surrogate model deems more likely to be good-performing than poor-performing. The ratio is monotonically related to the expected improvement acquisition function with threshold $\beta$.  [^2] Note that unlike in GP-based BO, we cannot set $\beta$ to the minimum observed value, as this would imply $\gamma=0$ and yield an empty set for fitting $l_{\beta}$. In practice, to approximate the maximization of $\alpha$, TPE samples a set of candidate configurations from $l_{\beta}$ and selects the one that maximizes the ratio $l_{\beta}(\lambda)/g_{\beta}(\lambda)$. This generative process of sampling $\lambda$ from $l_{\beta}$ allows TPE to flexibly accommodate conditional or hierarchical search spaces by employing appropriate factorizations of $l_{\beta}$ – for example, tree-structured models.
TPE is conceptually simple, naturally parallelizable, and handles complex, conditional search spaces well. Subsequent work has shown that the same acquisition objective can be approximated via probabilistic binary classification, which bypasses density estimation and leverages standard machine learning tools [Tiao:21].
4.2.6 Bayesian Optimization Vs. Random Search: An Example
@@
We illustrate the difference between model-based HPO and random search in a concrete example of tuning regularization parameters of a machine learning model to prevent overfitting. We tuned the L1 and L2 regularization terms, respectively alpha and lambda, of the XGBoost algorithm [chen2016xgboost] on the UCI direct marketing binary classification dataset from Kaggle. The goal was to minimize the AUC. Figure 4.2 presents a comparison between random search and BO (model-based), which models the objective function through a GP and uses the expected improvement as the acquisition function.
@@
4.3 Other Approaches to Model-Based HPO
Not all model-based methods follow a Bayesian optimization framework. ilievski2017efficient use a deterministic radial basis function model for $\hat{f}$ which they interpolate on observations of the response function. hazan2017hyperparameter focus on Boolean hyperparameters and use a sparse Fourier basis model for $\hat{f}$, in conjunction with heuristics to estimate the most sensitive hyperparameter entries. They then use a base global optimizer to perform HPO on a restriction of the original hyperparameter space. The works by lorraine2018stochastic and mackay2019self, related to gradient-based hyperparameter optimization (covered in Chapter 7), can also be interpreted as attempts to learn a surrogate of the entire learning algorithm (applied to a specific dataset), i.e., learning a mapping $\hat{f}\mathrel{\mathop{\ordinarycolon}}\Lambda\to\mathcal{H}$ that maps hyperparameters to trained models (cf. Eq. (2.1)).
Other approaches blend meta- and transfer-learning to model-based HPO by training a surrogate model on multiple datasets and tasks. In particular, [feurer2015initializing] extracts meta-features from multiple datasets, including descriptive statistics (number of instance, classes, mean, variance and higher moments) and “landmarks” (i.e. the performances of a few quick-to-evaluate learning algorithms on the dataset). They then use this information to train a surrogate model that initializes a BO process, thereby leveraging prior knowledge to accelerate optimization. fusi2018probabilistic follow instead a probabilistic matrix factorization approach akin to collaborative filtering in recommender systems, modeling the performance of configurations on datasets – therefore, users and items are “mapped to configurations and datasets). The learned probabilistic representations are then used to guide the search of next configurations to try via the expected improvement acquisition function.